Cost of Goods Sold Sexton Corp. has current liabilities of $263,000, a quick ratio of .75, inventory turnover of 10.35, and a current ratio of 1.25. What is the cost of goods sold for the company?
Answer #1

Current Assets = Current Ratio * Current liabilities = 1.25 * 263,000 = 328,750

Inventory = Current Assets - 0.75* Current Liabilities = 328,750 - 0.75* 263,000 = 131,500

COGS = Inventory Turnover * Inventory = 10.35 * 131500 = 1,361,025

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well